At a bottling company, machine A fills a bottle with spring water and machine B accepts the bottle only if the number of fluid ounces is between 11 1/8 and 12 1/8 If machine B accepts a bottle containing fluid ounces, which of the following describes all possible value of n ?
(A) |n-12| = 1/8
(B) |n+12| = 1/8
(C) |n-12| < 1/8
(D) |n+12| < 1/8
(E) |n-12| > 1/8

Answer by stanbon(75887)

At a bottling company, machine A fills a bottle with spring water and machine B accepts the bottle only if the number of fluid ounces is between 11 1/8 and 12 1/8 If machine B accepts a bottle containing fluid ounces, which of the following describes all possible value of n ?
----------------
Comment: Based on the 5 answers you listed, your posted information
is incorrect. The bottles should be filled between 11 7/8 and 12 1/8
fluid ounces.
----------------------------
1st: Find the middle value between 11 7/8 and 12 1/8
Ans: 12
----
2nd: Find the distance from the middle to one of the end points
Ans: 1/8
---
3rd: Since you want the difference to be "less than" the solution
is |n-12|< 1/8
======================
Cheers,
Stan H.
======================

(A) |n-12| = 1/8
(B) |n+12| = 1/8
(C) |n-12| < 1/8
(D) |n+12| < 1/8
(E) |n-12| > 1/8